Two particles A and B are released from rest, having equal charges of -5 x10-4 and initially separated from each other by 0.25m. The masses of particles A and B are 5.25x10-6 kg and 3.75x10-6 kg respectively. Find the magnitude of the initial accelerations of particles A and B.


1
Expert's answer
2022-02-18T08:22:45-0500

Given:

"q_1=q_2=\rm -5 *10^{-4}\: C"

"m_1=\rm 5.25*10^{-6}\: kg"

"m_2=\rm 3.75*10^{-6}\: kg"

"r=0.25\:\rm m"


The Coulomb's law says


"F=k\frac{|q_1||q_2|}{r^2}""F=9*10^9\frac{(5.0*10^{-4})^2}{0.25^2}=36000 \:\rm N"


The acceleration of particles

"a_1=\frac{F}{m_1}=\frac{36000}{5.25*10^{-6}}=6.9*10^9\:\rm m/s^2"

"a_2=\frac{F}{m_2}=\frac{36000}{3.75*10^{-6}}=9.6*10^9\:\rm m/s^2"